So my question is all about factoring.
It says to simplify all square roots and answers should not be decimals.
Factor or solve the following:
2x^3-3x^2-50x+75

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)
Answer by ikleyn(52800)

.

Factor it as follows


    2x^3 - 3x^2 - 50x + 75 = apply grouping first = (2x^3 - 3x^2) - (50x - 75) = 

                           = x^2*(2x-3) - 25*(2x-3) = (x^2-25)*(2x-3) = (x+5)*(x-5)*(2x-3).

Factoring is complete.